# Augmented Physics: A Machine Learning-Powered Tool for Creating Interactive Physics Simulations from Static Diagrams

### v1

@@ Abstract.
We introduce Augmented Physics, a machine learning-powered tool designed for creating interactive physics simulations from static textbook diagrams. Leveraging computer vision techniques, such as Segment Anything and OpenCV, our web-based system enables users to semi-automatically extract diagrams from physics textbooks and then generate interactive simulations based on the extracted content. These interactive diagrams are seamlessly integrated into scanned textbook pages, facilitating interactive and personalized learning experiences across various physics concepts, including gravity, optics, circuits, and kinematics. Drawing on an elicitation study with seven physics instructors, we explore four key augmentation techniques: 1) augmented experiments, 2) animated diagrams, 3) bi-directional manipulatives, and 4) parameter visualization. We evaluate our system through technical evaluation, a usability study (N=12), and expert interviews (N=12). The study findings suggest that our system can facilitate more engaging and personalized learning experiences in physics education.
@@ 1. Introduction
However, creating these interactive simulations is time-consuming and requires significant effort in programming. Given that most educators as well as students lack such technical skills, they need to rely on readily available ”off-the-shelf” simulations found online. While these generic simulations can be useful, they sometimes fail to match the exact needs and context of the students’ learning materials. For example, our formative study revealed that students frequently struggle to find external simulators that precisely align with textbook content. Moreover, they have to switch back and forth between the website and textbooks to contextualize the online example with their own learning materials. Due to this tedious and distracting process, dynamic simulations are not effectively utilized in current educational settings, despite their potential benefits.
@@
In this paper, we propose Augmented Physics, a novel approach to creating interactive physics simulations by augmenting static textbook diagrams. By leveraging advanced computer vision techniques like Segment-Anything (Kirillov et al., 2023) and OpenCV, users can semi-automatically extract diagrams from scanned textbook pages, and then generate interactive simulations based on the extracted content. Our system supports various types of simulations, such as Newtonian motion, optics, circuits, and repeated animation (Figure 1). Through a simple authoring process, users can select which objects to segment, manipulate segmented objects, and change parameter values to interact with simulation results. Furthermore, these interactive visual outputs are seamlessly overlaid onto the scanned PDF through a web-based interface, allowing students to learn, experiment, and play with their textbooks, without needing to search for external materials (Perkins et al., 2006) or create simulations from scratch (Scott and Davis, 2013; Cheema and LaViola, 2012).
@@
The idea of creating interactive explanations from static documents is not new (Masson et al., 2023; Chulpongsatorn et al., 2023), but this paper contributes in three key ways. First, we contribute a novel image-to-simulation pipeline. Existing works like Charagraph (Masson et al., 2023) and Augmented Math (Chulpongsatorn et al., 2023) mostly focus on text-to-text or text-to-graph pipelines using standard OCR or simple image recognition, but such pipelines do not suffice for physics simulations, which require more image-centric approach. Therefore, we develop a pipeline to segment objects, recognize images, convert them to simulation-ready objects, and integrate them into the scanned diagrams. To the best of our knowledge, our work is the first to explore and demonstrate this image-based physics simulator generation.
@@
Second, we contribute to the design space of augmented physics simulation tools. To design our system, we conducted a formative elicitation study, asking seven physics instructors about how they would augment a physics textbook. Based on the results, we identified four key augmentation strategies: 1) augmented experiments, 2) animated diagrams, 3) bi-directional manipulatives, and 4) parameter visualization.
@@
Third, we contribute insights from three evaluations: a technical evaluation, a preliminary usability study (N=12), and expert interviews with physics instructors (N=12). For technical evaluations, we assessed performance and accuracy across 200 diagrams from five different physics textbooks. Our technical evaluation results indicate that our pipeline varies based on the types of diagrams, such as kinematics (64%), optics (44%), circuits (15%), and animation (66%), based on 200 diagrams across six physics textbooks. Through the user study and expert interviews, we qualitatively compare our approach to existing learning practices, such as handouts, videos, and existing interactive websites to explore how our tool could meet their needs and fit into current practices. Their feedback suggests that our system complements, rather than replaces, existing learning materials such as videos and online simulators. While well-developed existing materials might better work for prepared topics, our tool serves on-demand and personalized learning assistance tailored to specific contexts, which is not well-supported by current learning practices. Based on their feedback and insights, we discuss ways to expand our proposed approach beyond the current proof-of-concept prototype for future deployment.
@@
Augmented Physics111The code and online demos are publicly available as open source at [anonymized]., a tool for creating interactive simulations by extracting and animating static physics diagrams.
@@ 2.1. Physics Simulation for Learning
However, these existing physics simulation tools are often limited to pre-programmed and off-the-shelf simulations, which sometimes fail to meet the specific needs and challenges students face. To address this limitation, HCI researchers have explored authoring tools that allow users to create personalized physics simulations on demand. Tools like PhysInk (Scott and Davis, 2013), PhysicsBook (Cheema and LaViola, 2012), MathPad2 (LaViola and Zeleznik, 2007), and ChalkTalk (Perlin et al., 2018) enable users to sketch physics diagrams, which are then automatically transformed into interactive and animated graphics. Such authoring tools allow non-technical users like students and instructors to easily and quickly create physics simulations without programming skills. While these sketching tools hold great potential as classroom instruction tools for teachers, our formative study reveals that students often struggle to create effective simulations from scratch by themselves without clear guidance and instructions. To fill this gap, this paper explores an alternative approach to generating interactive simulations by animating existing static diagrams instead of sketching them from scratch.
@@ 2.2. Augmenting Existing Documents
Previous research has investigated how to make static explanations more dynamic and interactive. For instance, Victor introduced the concept of Explorable Explanations (Victor, 2011a), demonstrating various interactive explanations for scientific reading (Victor, 2005, 2011b, 2013a). Such interactive explanations enhance readers’ understanding by highlighting the relationship between texts and data (Chen and Xia, 2022; Latif et al., 2021; Kong et al., 2014), allowing in-situ exploration through multiverse analysis (Dragicevic et al., 2019) among others (Hohman et al., 2020). However, a key limitation of interactive documents is the inherent need for programming, which requires substantial time and cost to create them (Head et al., 2022). While tools like Tangle (Victor, 2013b), Idyll Studio (Conlen et al., 2021), and Data Theater (Lau and Guo, 2020) aim to lower this barrier, they still need programming, leaving existing static documents unusable for interactive explanations.
To address this problem, researchers have developed methods to augment existing documents, rather than programming them from scratch. Prior works have investigated tools to semi-automatically generate summaries (e.g., Marvista (Chen et al., 2022)), references (e.g., HoloDoc (Li et al., 2019)), highlights (e.g., ScentHighlights (Chi et al., 2005), Scim (Fok et al., 2023), Kim et al. (Kim et al., 2018)), annotations (e.g., Threddy (Kang et al., 2022), Contextifier (Hullman et al., 2013), textSketch (Subramonyam et al., 2020), DuallyNoted (Qian et al., 2022)), and visualizations (e.g., Elastic Documents (Badam et al., 2018), Jessica et al. (Hullman et al., 2018)) by augmenting existing documents. Most closely related to our work, Charagraph (Masson et al., 2023) and Augmented Math (Chulpongsatorn et al., 2023) explore the semi-automatic generation of interactive charts and graphs by extracting text from static documents. However, these text-based extraction methods fall short for physics simulations, which heavily rely on visual representation, rather than textual input. Thus, our paper contributes a new pipeline to generate interactive explanations directly from images, by leveraging object segmentation, image recognition, and physics simulator generation.
@@ 2.3. Tools for Authoring Interactive Diagrams
Previous research has explored end-user authoring tools for creating dynamic and interactive diagrams for various applications, including technical illustrations (Zhu et al., 2011), scientific explanation (Sarracino et al., 2017), and artistic animation (Xing et al., 2016). In the educational domain, many online tools (GeoGebra, [n.d.]; Walsh, [n.d.]) and research prototypes (Saquib et al., 2021; LaViola and Zeleznik, 2007) allow for interactive authoring and animation. These tools enable users to create animation through sketch-based interactions (Willett et al., 2018; Davis et al., 2008; Kazi et al., 2014b; Kazi et al., 2016; Kazi et al., 2014a) and tangible demonstrations (Barnes et al., 2008). Such authoring techniques have been demonstrated to be versatile and adaptable across various domains, including 3D animations (Ma et al., 2022), video augmentation (Hashim et al., 2023; Xia et al., 2023), and motion graphics videos (Jahanlou and Chilana, 2022). Although these methods have significantly enhanced the creative authoring of dynamic visuals, they may not be ideally tailored for explanatory content in physics textbooks. In these contexts, animated objects must adhere to specific physical behaviors, making the general techniques potentially less effective.
@@ 3.3. Elicited Augmentation Strategies
Bi-Directional Manipulatives
@@ 4.1. Overview
In this section, we introduce Augmented Physics, a machine learning-powered authoring tool designed to enable non-technical users to create interactive physics simulations from static diagrams. Our web-based tool facilitates users, including students and instructors, in semi-automatically extracting diagrams from physics textbooks and generating simulations that seamlessly integrate with scanned textbook pages. Our research primarily focuses on basic physics concepts taught in high schools across the United States, such as Newtonian motion, optics, and electrical circuits. Although more advanced topics like quantum mechanics are beyond our current scope, our adaptable animated diagrams technique allows users to create animated illustrations for these concepts as well. Moreover, we have made our system open-source, including the machine learning pipeline and browser-based simulators, to encourage the HCI community to further develop our prototype and methods.
@@ 4.2. Authoring Workflow
Upon importing the diagram, the system asks the user to select a type of simulation from the available options. Users have the choice among three specific simulation categories: kinematics, optics, and circuits. There is also an option for animation, catering to scenarios that do not necessitate a particular type of simulation.
@@
Once users have segmented images and assigned roles, the system proceeds to generate the simulation by converting the segmented images into polygons with appropriate properties for the physics simulation. For example, the skier and slope shown in Figure 7 are precisely replicated to create polygons for dynamic and static objects, respectively. This approach ensures the simulation integrates seamlessly with the original diagram, achieving alignment in both shape and position within the image. After completing these steps, the simulation is ready to be launched. Users can start it by clicking the run button 7 or by interacting with the rendered polygons to witness dynamic visual outputs (Figure 8). They can click on the simulated objects and optionally change parameters.
@@
Users have the flexibility to adjust parameters within the simulation. Depending on their roles, different objects come with various parameters, such as mass for dynamic objects, friction for static objects, and force constants for springs. Moreover, the system can recognize parameter values within text or images, enabling users to manipulate numerical values on the page. For example, in electrical circuit simulations, users can modify values like those of resistors and batteries to dynamically change the simulation results. Additionally, users can link numerical values from the text to specific properties of objects in the simulation by simply binding these values.
@@ 4.3. Supported Augmentation Features
We have developed the following four augmentation techniques: 1) augmented experiments, 2) animated diagrams, 3) bi-directional manipulatives, and 4) parameter visualization. These features are tailored to support a wide range of simulated experiments that educators wish to create for their students.
@@ 4.3.1. Augmented Experiments
As previously stated, our system supports three simulation categories: 1) kinematics, covering gravity, Newtonian motion, springs, and pendulums; 2) optics, focusing on lenses, light reflection, and image formation through mirroring; and 3) electrical circuits, dedicated to simulating current flow in electronic circuits.
@@ 4.3.2. Animated Diagrams
4.3.3. Bi-Directional Manipulatives
Bi-directional manipulatives enable authors to link parameter values from the text to the associated simulation. This feature allows students to adjust these values directly within the text and observe the changes in real time. Initially, the system identifies and highlights all numbers within the provided image for the author. Then, the author can select a specific numerical value from the text and assign it a simulation property through a dropdown menu that displays all available properties. For example, Figure 11 illustrates how the user binds the value in the text to the compression property, enabling the system to use this value to simulate the scenario by changing the spring’s compression.
@@ 4.4. Implementation
Our system comprises two main components: a backend computer vision pipeline module using Python and a frontend web interface developed with React.js. The computer vision module integrates Segment-Anything (Kirillov et al., 2023), a widely utilized image segmentation model, alongside custom-developed line and contour detection algorithms through OpenCV. Communication between the frontend and backend is facilitated via the Firebase real-time database, allowing for the processing of images based on provided input coordinates. The results, including extracted images, lines, or points, are subsequently communicated back through Firebase. Moreover, we compute the bounding boxes and X and Y coordinates of extracted image segments and transmit this data to the frontend. For text recognition and the extraction of numerical values, we utilize Google’s Cloud Vision API. For our prototype and technical evaluation, we used Google Chrome on MacBook Air 14 inch 2022 (M2 with 10-Core Integrated GPU and 16GB RAM) for the frontend and Google Colab (CPU: Intel Xeon 4 cores, GPU: Nvidia T4, RAM: 50GB) for the backend.

### v2

# Augmented Physics: Creating Interactive and Embedded Physics Simulations from Static Textbook Diagrams

@@ Abstract.
We introduce Augmented Physics, a machine learning-integrated authoring tool designed for creating embedded interactive physics simulations from static textbook diagrams. Leveraging recent advancements in computer vision, such as Segment Anything and Multi-modal LLMs, our web-based system enables users to semi-automatically extract diagrams from physics textbooks and generate interactive simulations based on the extracted content. These interactive diagrams are seamlessly integrated into scanned textbook pages, facilitating interactive and personalized learning experiences across various physics concepts, such as optics, circuits, and kinematics. Drawing from an elicitation study with seven physics instructors, we explore four key augmentation strategies: 1) augmented experiments, 2) animated diagrams, 3) bi-directional binding, and 4) parameter visualization. We evaluate our system through technical evaluation, a usability study (N=12), and expert interviews (N=12). Study findings suggest that our system can facilitate more engaging and personalized learning experiences in physics education.
Teaser figures with various augmented physics features like augmented experiments and bi-directional manipulatives.
@@ 1. Introduction
However, creating these interactive simulations is time-consuming and requires significant effort in programming. Given that most educators, as well as students, lack such technical skills, they need to rely on readily available ”off-the-shelf” simulations found online. While these generic simulations can be useful, they sometimes fail to match the exact needs and context of the students’ learning materials. For example, our formative study revealed that students frequently struggle to find external simulators that precisely align with textbook content. Moreover, they have to switch back and forth between the website and textbooks to contextualize the online example with their own learning materials. Due to this tedious and distracting process, dynamic simulations are not effectively utilized in current educational settings despite their potential benefits.
@@
In this paper, we propose Augmented Physics, a novel approach to creating interactive physics simulations by extracting and augmenting content from static textbook diagrams. By leveraging advanced computer vision techniques like Segment-Anything (Kirillov et al., 2023) and Multi-modal LLMs, teachers and educators can semi-automatically extract diagrams from textbook pages and generate interactive simulations based on the extracted content. Our system supports various types of simulations, such as Newtonian motion, optics, circuits, and looping animation (Figure 1). Through a simple authoring process, users can select specific objects in the diagram to segment, manipulate these segmented objects, and adjust parameter values to dynamically interact with simulation results. Furthermore, these interactive visual outputs are seamlessly overlaid onto the textbook PDF through a web-based interface, allowing students to learn, experiment, and play with their textbooks without needing to search for external materials (Perkins et al., 2006) or create simulations from scratch (Scott and Davis, 2013; Cheema and LaViola, 2012).
@@
The idea of creating interactive explanations from static documents is not new (Masson et al., 2023; Chulpongsatorn et al., 2023), but this paper contributes in three key ways. First, we contribute a novel image-to-simulation pipeline. Existing works like Charagraph (Masson et al., 2023) and Augmented Math (Chulpongsatorn et al., 2023) mainly focus on text-to-text or text-to-graph pipelines using standard OCR or simple image boundary detection, but such pipelines do not suffice for physics diagrams and simulations, which require a more image-centric approach. Therefore, we develop a pipeline to segment diagrams, recognize images, convert them to simulation-ready objects, and integrate them into the textbook diagrams. To the best of our knowledge, our work is the first to explore and demonstrate this image-based physics simulation generation.
@@
Second, we contribute to the design space of augmented physics simulation tools. To design our system, we conducted a formative elicitation study, asking seven physics instructors about how they would augment a physics textbook. Based on the results, we identified four key augmentation strategies: 1) augmented experiments, 2) animated diagrams, 3) bi-directional binding, and 4) parameter visualization.
Third, we contribute insights from three evaluations: a technical evaluation, a preliminary usability study (N=12), and expert interviews with physics instructors (N=12).
@@
Our technical evaluation results indicate that our pipeline varies based on the types of diagrams, such as kinematics (64%), optics (44%), circuits (40%), and animation (66%), drawn from 200 diagrams across six physics textbooks (these scores reflect the simulation pipeline operating without any modifications, but they are higher when small adjustments are made). Through the user study and expert interviews, we qualitatively compare our approach to existing learning practices, such as handouts, videos, and existing interactive websites, to explore how our tool could meet their needs and fit into current educational practices. Their feedback suggests that our system complements, rather than replaces, existing learning materials such as videos and online simulators. While well-developed existing materials might better work for prepared topics, our tool serves as a way for educators to create on-demand and personalized learning material tailored to specific contexts, which is not well-supported by current practices. Based on their feedback and insights, we discuss ways to expand our proposed approach beyond the current proof-of-concept prototype for future deployment.
@@
Augmented Physics111The code and demos are available at https://adigunturu.github.io/AugmentedPhysics/., a tool for creating interactive simulations by extracting and animating static physics diagrams.
@@ 2.1. Physics Simulation for Learning
However, these existing physics simulation tools are often limited to pre-programmed and off-the-shelf simulations, which sometimes fail to meet the specific needs and challenges students face. To address this limitation, HCI researchers have explored authoring tools that allow users to create personalized physics simulations on demand. Tools like PhysInk (Scott and Davis, 2013), PhysicsBook (Cheema and LaViola, 2012), MathPad2 (LaViola and Zeleznik, 2007), and ChalkTalk (Perlin et al., 2018) enable users to sketch physics diagrams, which are then automatically transformed into interactive and animated graphics. Such authoring tools allow non-technical users like students and instructors to easily and quickly create physics simulations without programming skills. While these tools hold great potential, they do not focus on simulating existing figures but on creating them from scratch, which can become tedious. Textbooks are primary teaching materials and contain rich and expressive diagrams for various concepts, which can be used to create personalized and situated learning experiences for students. Moreover, extracting textbook content also allows teachers to create higher-quality simulations that are identical and situated to student’s textbooks (which they can take home) compared to low-fidelity sketches. Our formative study reveals that students often need clear guidance and instructions to find related resources. To fill this gap, this paper explores an alternative approach to generating interactive simulations by animating existing static diagrams instead of sketching them from scratch.
@@ 2.2. Augmenting Existing Documents
Previous research has investigated how to make static explanations more dynamic and interactive. For instance, Victor introduced the concept of Explorable Explanations (Victor, 2011a), demonstrating various interactive explanations for scientific reading (Victor, 2005, 2011b, 2013a). Such interactive explanations enhance readers’ understanding by highlighting the relationship between texts and data (Latif et al., 2021; Kong et al., 2014), allowing in-situ exploration through multiverse analysis (Dragicevic et al., 2019) among others (Hohman et al., 2020). However, a key limitation of interactive documents is the inherent need for programming, which requires substantial time and cost to create them (Head et al., 2022). While tools like Tangle (Victor, 2013b), Idyll Studio (Conlen et al., 2021), and Data Theater (Lau and Guo, 2020) aim to lower this barrier, they still need programming, leaving existing static documents unusable for interactive explanations.
To address this problem, researchers have developed methods to augment existing documents, rather than programming them from scratch. Prior works have investigated tools to semi-automatically generate summaries (e.g., Marvista (Chen et al., 2022)), references (e.g., HoloDoc (Li et al., 2019)), highlights (e.g., ScentHighlights (Chi et al., 2005), Scim (Fok et al., 2023), Kim et al. (Kim et al., 2018)), annotations (e.g., Threddy (Kang et al., 2022), Contextifier (Hullman et al., 2013), textSketch (Subramonyam et al., 2020), DuallyNoted (Qian et al., 2022)), and visualizations (e.g., Elastic Documents (Badam et al., 2018), Jessica et al. (Hullman et al., 2018)) by augmenting existing documents. Most closely related to our work, Charagraph (Masson et al., 2023) and Augmented Math (Chulpongsatorn et al., 2023) explore the semi-automatic generation of interactive charts and graphs by extracting text from static documents. While our work shares a similar motivation with some others: to make textbooks interactive and explorable, our goals and methods differ fundamentally. While other works focus on augmenting text content, our system focuses on making diagrams themselves explorable and animated by extracting and simulating individual components. This is crucial for Physics education, where diagrams represent dynamic systems and processes that change over time. High school physics concepts rely on visualizing motion, and our augmentation strategies enable users to animate individual components in the image and make them interact with each other to craft engaging and explorable experiences.
@@ 2.3. Tools for Authoring Interactive Diagrams
Previous research has explored end-user authoring tools for creating dynamic and interactive diagrams for various applications, including technical illustrations (Zhu et al., 2011), scientific explanation (Sarracino et al., 2017), and artistic animation (Xing et al., 2016). In the educational domain, many online tools (GeoGebra, [n.d.]; Walsh, [n.d.]) and research prototypes (Saquib et al., 2021; LaViola and Zeleznik, 2007) allow for interactive authoring and animation. These tools enable users to create animation through sketch-based interactions (Willett et al., 2018; Davis et al., 2008; Kazi et al., 2014b; Kazi et al., 2016; Kazi et al., 2014a) and tangible demonstrations (Barnes et al., 2008). Such authoring techniques have been demonstrated to be versatile and adaptable across various domains, including 3D animations (Ma et al., 2022), video augmentation (Hashim et al., 2023; Xia et al., 2023a), and motion graphics videos (Jahanlou and Chilana, 2022). Although these methods have significantly enhanced the creative authoring of dynamic visuals, they may not be ideally tailored for explanatory content in physics textbooks. In these contexts, animated objects must adhere to specific physical behaviors, making the general techniques potentially less effective.
@@ 3.3. Elicited Augmentation Strategies
4 sketches depicting various examples of Augmented Experiments.
@@
4 sketches depicting various examples of Animated Diagrams.
Bi-Directional Binding
@@
4 sketches depicting various examples of Bi-Directional Binding.
@@
4 sketches depicting various examples of Parameter Visualization.
@@ 4.1. Overview
In this section, we introduce Augmented Physics, a machine learning-integrated authoring tool designed to enable non-technical users to create interactive physics simulations from static diagrams. Our web-based tool facilitates users, including students and instructors, in semi-automatically extracting diagrams from physics textbooks and generating simulations that seamlessly integrate with scanned textbook pages. Our research primarily focuses on basic physics concepts taught in high schools across the United States, such as Newtonian motion, optics, and electric circuits. Although more advanced topics like quantum mechanics are beyond our current scope, our adaptable animated diagrams technique allows users to create animated illustrations for these concepts as well. Moreover, we have made our system open-source, including the machine learning pipeline and browser-based simulators, to encourage the HCI community to further develop our prototype and methods.
3 images of an optics diagram being simulated with interactivity.
@@ 4.2. Authoring Workflow
Upon importing the diagram, the system asks the user to select a type of simulation from the available options. The system first automatically recommends a simulation type. Users have the choice among three specific simulation categories: kinematics, optics, and circuits. There is also an option for animation, catering to scenarios that do not necessitate a particular type of simulation.
@@
3 images of a kinematic diagram being simulated as an augmented experiment.
@@
Once users have segmented images and assigned roles, the system proceeds to generate the simulation by converting the segmented images into polygons with appropriate properties for the physics simulation. For example, the skier and slope shown in Figure 7 are precisely replicated to create polygons for dynamic and static objects, respectively. This approach ensures the simulation integrates seamlessly with the original diagram, achieving alignment in both shape and position within the image. After completing these steps, the simulation is ready to be launched. Users can start it by clicking the run button Figure 7 or by interacting with the rendered polygons to witness dynamic visual outputs (Figure 8). They can click on the simulated objects and optionally change parameters.
@@
Users have the flexibility to adjust parameters within the simulation. Depending on their roles, different objects come with various parameters, such as mass for dynamic objects, friction for static objects, and force constants for springs. Moreover, the system can recognize parameter values within text or images, enabling users to manipulate numerical values on the page. For example, in electrical circuit simulations, users can modify values like those of resistors and batteries to dynamically change the simulation results. Additionally, the system automatically links numerical values from the text to specific properties of objects in the simulation, which the user can edit.
Another 3 images of an optics diagram being simulated with interactivity.
@@ 4.3. Supported Augmentation Features
We have developed the following four augmentation techniques: 1) augmented experiments, 2) animated diagrams, 3) bi-directional binding, and 4) parameter visualization. These features are tailored to support a wide range of simulated experiments that educators wish to create for their students.
@@ 4.3.1. Augmented Experiments
Images of circuits being simulated with animations.
As previously stated, our system supports three simulation categories: 1) kinematics, covering gravity, Newtonian motion, springs, and pendulums; 2) optics, focusing on lenses, light propagation, and image formation through mirroring; and 3) electric circuits, dedicated to simulating current flow in electronic circuits.
@@ 4.3.2. Animated Diagrams
Images of animated diagrams in a physics textbook.
4.3.3. Bi-directional Binding
Bi-directional Binding enable authors to link parameter values from the text to the associated simulation. This feature allows students to adjust these values directly within the text and observe the changes in real time. Initially, the system identifies and highlights all numbers within the provided image for the author. Then, the author can select a specific numerical value from the text and assign it a simulation property through a dropdown menu that displays all available properties. For example, Figure 11 illustrates how the user binds the value in the text to the compression property, enabling the system to use this value to simulate the scenario by changing the spring’s compression.
Images of physics textbook being simulated with bi-directional binding.
@@ 4.3.4. Parameter Visualization
Images of a pendulum being simulated in-situ overlaid on the textbook.
@@ 4.4. Implementation
Our system comprises two main components: a backend computer vision pipeline module using Python and a frontend web interface developed with React.js. The computer vision module integrates Segment-Anything (Kirillov et al., 2023), a widely utilized image segmentation model, alongside custom-developed line and contour detection algorithms through OpenCV. Communication between the frontend and backend is facilitated via the Firebase real-time database, allowing for the processing of images based on provided input coordinates. The results, including extracted images, lines, or points, are subsequently communicated back through Firebase. Moreover, we compute the bounding boxes and X and Y coordinates of extracted image segments and transmit this data to the frontend. For text recognition and the extraction of numerical values, we utilize Google’s Cloud Vision API. We then send the page text and extracted object data in JSON format to an LLM (GPT-4), which is used to recommend simulation type and automatically set the parameters of the simulations based on text. For our prototype and technical evaluation, we used Google Chrome on MacBook Air 14 inch 2022 (M2 with 10-Core Integrated GPU and 16GB RAM) for the frontend and Google Colab (CPU: Intel Xeon 4 cores, GPU: Nvidia T4, RAM: 50GB) for the backend.